# Asymptotic theory of in-context learning by linear attention

### v2

@@ I.1 Related Work
Many theoretical studies of ICL have centered on the idea that Transformers learn a particular algorithm during pretraining, which is then flexibly deployed to solve in-context tasks. In broad strokes, papers from this program of research often consider a particular algorithm for solving an in-context task, prove that Transformers can approximately implement this algorithm, and then empirically compare the ICL performance of a pre-trained Transformer to the performance of that algorithm [22, 23, 24, 25, 26, 27, 28, 29]. A clear consensus on which algorithm underlies ICL of linear regression in full transformers has yet to emerge [22, 23, 24, 25, 26, 27, 28, 29]. Within this line of research, closest to our work are a series of papers that consider ICL of linear regression by simplified Transformers using linear, rather than softmax, attention modules [29, 28, 25, 30, 27, 31, 32]. Zhang, Frei, and Bartlett [29] studied these models in the limit of infinite pretraining dataset size (i.e., the population risk limit), and show that their performance on in-context linear regression nearly matches that of the Bayes-optimal estimator for the ICL task. However, they found that linear Transformers are not robust to shifts in the within-context covariate distribution. Zhang, Wu, and Bartlett [28] then showed that any optimizer of the within-context risk for a linear Transformer solves the ICL task with an approximation to one step of gradient descent from a learnable initialization, and that the resulting estimator can saturate the Bayes error for tasks with a Gaussian prior and non-zero mean. As we will discuss in Section II, our reduction of the linear attention module is inspired in part by these works. In very recent work, Duraisamy [32] has studied the fine-sample risk of in-context linear regression with a single step of gradient descent, without directly analyzing Transformers. An et al. [25] and Wu et al. [30] investigated how linear Transformers adapt to limited pretraining data and context length, again showing that in certain cases nearly-optimal error is achievable. Like these studies, our work considers linear attention; but our analysis, with its asymptotically sharp predictions, allows us to elucidate the exact dependence of ICL performance on various hyperparameters and to pinpoint when and how the transition from memorization to ICL of linear regression occurs. In closing, we highlight the work of Reddy [21], who analyzed the emergence of in-context classification through a phenomenological model.
@@ II.2 Linear self-attention
where 00 in the lower-right corner is a token that prompts the missing value $y_{\ell+1}$ to be predicted.
@@
The output $A$ is a matrix while our goal is to predict a scalar, $y_{\ell+1}$. Following the choice of positional encoding in (4), we will take $A_{d+1,\ell+1}$, the element of $A$ corresponding to the 00 prompt, as the prediction for $y_{\ell+1}$:
@@ II.3 Pretraining data
The model is pretrained on $n$ sample sequences, where the $\mu$th sample is a collection of $\ell+1$ vector-scalar pairs $\{x_{i}^{\mu}\in\mathbb{R}^{d},y_{i}^{\mu}\in\mathbb{R}\}_{i=1}^{\ell+1}$ related by the approximate linear mapping in (1): $y_{i}^{\mu}=\langle x_{i}^{\mu},w^{\mu}\rangle+\epsilon_{i}^{\mu}$. Here, $w^{\mu}$ denotes the task vector associated with the $\mu$th sample. We make the following statistical assumptions:
@@
    For $1\leq\mu\leq n$, the task vector $w^{\mu}$ associated with the $\mu$th sample context is uniformly sampled from $\Omega_{k}$. Note that the variable $k$ controls the task diversity in the pretraining dataset. Importantly, $k$ can be less than $n$, in which case the same task vector from $\Omega_{k}$ will be repeated multiple times.
@@ II.4 Parameter reduction
[…]
@@
We note that [29] provides an analysis of the population risk (whereas we focus on empirical risk) for a related reduced model in which both $v_{21}$ and $m_{21}$ are set to 00. Consequently, the predictors considered in [29] slightly differ from ours in (16)–(18) by an additive term (due to $m_{21}$). The authors of [29] justify this reduced model through an optimization argument: if these parameters $v_{21}$ and $m_{21}$ are initialized to zero, they remain zero under gradient descent optimization of the population risk.
@@ II.5 Model pretraining
$\Gamma^{\ast}$ $=\underset{\Gamma}{\arg\,\min}\,\sum_{\mu=1}^{n}\left(y_{\ell+1}^{\mu}-\langle\Gamma,H_{Z^{\mu}}\rangle\right)^{2}+\frac{n}{d}\lambda\lVert\Gamma\rVert_{\mathrm{F}}^{2},$ (20)
where $\lambda>0$ is a regularization parameter, and $H_{Z^{\mu}}$ refers to the input matrix (17) populated with the $\mu$th sample sequence. The factor $n/d$ in front of $\lambda$ makes sure that, when we take the $d\to\infty$ or $n\to\infty$ limits later, there is still a meaningful ridge regularization when $\lambda>0$. The solution to the optimization problem in (20) can be expressed explicitly as
$\text{vec}(\Gamma^{\ast})=\left(\frac{n}{d}\lambda I+\sum_{\mu=1}^{n}\text{vec}(H_{Z^{\mu}})\text{vec}(H_{Z^{\mu}})^{\top}\right)^{-1}\sum_{\mu=1}^{n}y_{\ell+1}^{\mu}\text{vec}(H_{Z^{\mu}})$ (21)
@@ III.2 Learning curves for ICL of linear regression by a linear attention module
where $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ and $e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)$ are two deterministic functions of the parameters $\tau$, $\alpha$, $\kappa$, $\rho$ and $\lambda$. The exact expressions of these two functions can be found in Section SI.5.2 and Section SI.5.3 of the SI, respectively. For simplicity, we only present in what follows the ridgeless limit (i.e., $\lambda\to 0^{+}$) of the asymptotic generalization errors.
@@
where $\mathcal{M}_{\kappa}(\cdot)$, defined in (SI.8.3), is a function related to the Stieltjes transform of the Marchenko-Pastur law. Then
@@
[…]
@@
We will discuss various implications of these equations in the next sections. We derived these results using techniques from random matrix theory. The full setup and technical details are presented in the Supplementary Information in Section SI.1 through Section SI.5. The computations involve analysis of the properties of the finite-sample optimal parameter matrix $\Gamma^{*}$ (see (21)). A key technical component of our analysis involves characterizing the spectral properties of the sample covariance matrix of $n=\Theta(d^{2})$ i.i.d. random vectors in dimension $\Theta(d^{2})$. Each of these vectors is constructed as the vectorized version of the matrix in (17). Related but simpler versions of this type of random matrices involving the tensor product of i.i.d. random vectors have been studied in recent work [39]. Some of our derivations are based on non-rigorous yet technically plausible heuristics. We support these predictions with numerical simulations and discuss in the Supplementary Information the steps required to achieve a fully rigorous proof.
@@ III.3 Sample-wise double-descent
((a)) Ridgeless IDG error against $\tau$
((b)) Ridgeless ICL error against $\tau$
((c)) Finite ridge ICL error against $\tau$
Figure 1: ICL performance as a function of $\tau$: theory (solid lines) vs simulations (dots). Plots of (a) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$, (b) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$, and (c) $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ against $\tau$. Simulated errors calculated by evaluating the corresponding test error on the corresponding optimized $\Gamma^{*}$. Parameters: $d=100$, $\rho=0.01$ for all; (a), (b) $\kappa=0.5$, (c) $\alpha=10,\kappa=\infty$. Averages and standard deviations are computed over 10 runs.
How large should $n$, the pretraining dataset size, be for the linear attention to succesfully learn the task in-context? In Figure 1, we plot our theoretical predictions for the ICL and IDG error as a function of $\tau=n/d^{2}$ and verify them with numerical simulations. Our results demonstrate that the quadratic scaling of sample size with input dimensions is indeed an appropriate regime where nontrivial learning phenomena can be observed.
As apparent in Figure 1, we find that the generalization error for both ICL and IDG tasks are not monotonic in the number of samples. In the ridgeless limit, both ICL and IDG errors diverge at $\tau=1$, with the leading order behavior in the $\tau\uparrow 1$ (respectively $\tau\downarrow 1$) limit given by $\frac{c_{1}}{1-\tau}$ (respectively $\frac{c_{2}}{\tau-1}$), where $c_{1}$ (respectively $c_{2}$) is a $\tau$-independent constant. This leads to a “double-descent” behavior [40, 38] in the number of samples. As in other models exhibiting double-descent [40, 38, 41], the location of the divergence is at the interpolation threshold: the number of parameters of the model (elements of $\Gamma$) is, to leading order in $d$, equal to $d^{2}$, which matches the number of pretraining samples at $\tau=1$. Further, we can investigate the effect of ridge regularisation on the steepness of the double descent, as illustrated in Figure 1(c) for the ICL task. As we would expect from other models exhibiting double-descent [40, 38, 41], increasing the regularization strength suppresses the peak in error around the interpolation threshold.
Finally, we note that if we take the limit of $\kappa\to\infty$ and $\alpha\to\infty$ in Result 1 (in either order), the ICL generalization error in the ridgeless case reduces to the generalization error of simple ridgeless interpolation with isotropic Gaussian covariates in $d^{2}$ dimensions [38, 41]:
$\lim_{\alpha\to\infty}\lim_{\kappa\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ $=\lim_{\kappa\to\infty}\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ (33)
=$=\begin{dcases}1-\tau+\frac{\rho}{1-\tau}&\tau<1,\\ \frac{\rho\tau}{\tau-1}&\tau>1.\end{dcases}$ (34)
@@ III.4 ICL and IDG error curves can have non-monotonic dependence on context length
((a)) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ against $\alpha$; $\tau=0.5,\rho=0.5$
((b)) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ against $\alpha$; $\tau=0.5,\rho=0.1$
((c)) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ against $\alpha$; $\tau=20,\rho=0.5$
Figure 2: Error curves as functions of $\alpha$: theory (solid lines) vs simulations (dots). Plots of (a) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ and (b), (c) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ against $\alpha$. Hollowed markers, when plotted, indicate $\alpha^{\ast}$ value minimizing error if it exists. Parameters: $d=100$. Averages and standard deviations are computed over 20-100 runs.
((a)) Non-monotonicity of $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$, $\tau=0.5$
((b)) Non-monotonicity of $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$, $\tau=0.5$
((c)) Non-monotonicity of $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$, $\tau=20$
Figure 3: Phase diagram of non-monotonicity in $\alpha$ of ridgeless IDG and ICL error, for fixed $\tau$ against task diversity $\kappa$ and label noise $\rho$. The colormap corresponds to the value $\alpha^{\ast}$ that minimizes error in $\alpha$, if it exists, at that particular $\kappa,\rho$ pair; grey is plotted if the error curve is monotonic at that $\kappa,\rho$ pair. The dashed vertical lines are plotted at $\kappa=\text{min}(\tau,1)$. We know from (35) and surrounding discussion that ICL error diverges in $\alpha$ for all $\kappa,\rho$ to the left of this line, and IDG error diverges on this line for $\tau<1$. Letters correspond to matching parameter values of the curves plotted in their respective Figure 2 subfigure.
How large should the context length be? In Figure 2, we plot our theoretical results verified with experiments. We observe that we have correctly identified the regime where in-context and in-weight learning appear: context length indeed scales linearly with input dimensions, as numerical simulations computed using finite $d$ fit the asymptotic error curves.
@@
While the functional form of $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ and $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ is too complex to study their minimizers analytically, we can investigate the non-monotonicity of IDG and ICL error in $\alpha$ numerically. This is done in Figure 3, where the blue-green colormap corresponds to the value $\alpha^{\ast}$, if it exists, that minimizes $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$ or $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$ at fixed $\tau,\kappa,\rho$. The grey color indicates cases where the corresponding function is monotonic in $\alpha$. The results reveal that the non-monotonicity of the errors with respect to context length $\alpha$ depends in a nontrivial manner on both the task diversity $\kappa$ and the noise-to-signal ratio $\rho$.
@@
$\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ =$=\begin{dcases}\infty&\kappa\leq\text{min}(\tau,1),\\ 1-\tau+\rho+\frac{\rho\kappa\tau}{(\kappa-\tau)(1-\tau)}&\tau<1,\kappa>\tau,\\ \rho+\frac{\rho\kappa}{(\kappa-1)(\tau-1)}&\tau>1,\kappa>1.\end{dcases}$ (35)
Note the explicit divergence of ICL error for $\kappa\leq\text{min}(\tau,1)$ (independent of $\rho$). Similar investigation of Result 2 under the limit $\alpha\to\infty$ reveals an explicit divergence of IDG error for $\kappa=\tau$. The divergence in $\alpha$ of ICL error is a peculiarity of the ridgeless limit of errors we have taken in Results 1 and 2, and is further discussed in Section SI.5.4 of the Supplementary Information.

### v3

@@ I.1 Related Work
Many theoretical studies of ICL have centered on the idea that Transformers learn a particular algorithm during pretraining, which is then flexibly deployed to solve in-context tasks. In broad strokes, papers from this program of research often consider a particular algorithm for solving an in-context task, prove that Transformers can approximately implement this algorithm, and then empirically compare the ICL performance of a pre-trained Transformer to the performance of that algorithm [22, 23, 24, 25, 26, 27, 28, 29]. A clear consensus on which algorithm underlies ICL of linear regression in full transformers has yet to emerge [22, 23, 24, 25, 26, 27, 28, 29]. Within this line of research, closest to our work are a series of papers that consider ICL of linear regression by simplified Transformers using linear, rather than softmax, attention modules [29, 28, 25, 30, 27, 31, 32]. Zhang, Frei, and Bartlett [29] studied these models in the limit of infinite pretraining dataset size (i.e., the population risk limit), and show that their performance on in-context linear regression nearly matches that of the Bayes-optimal estimator for the ICL task. However, they found that linear Transformers are not robust to shifts in the within-context covariate distribution. Zhang, Wu, and Bartlett [28] then showed that any optimizer of the within-context risk for a linear Transformer solves the ICL task with an approximation to one step of gradient descent from a learnable initialization, and that the resulting estimator can saturate the Bayes error for tasks with a Gaussian prior and non-zero mean. As we will discuss in Section˜II, our reduction of the linear attention module is inspired in part by these works. In very recent work, Duraisamy [32] has studied the fine-sample risk of in-context linear regression with a single step of gradient descent, without directly analyzing Transformers. An et al. [25] and Wu et al. [30] investigated how linear Transformers adapt to limited pretraining data and context length, again showing that in certain cases nearly-optimal error is achievable. Like these studies, our work considers linear attention; but our analysis, with its asymptotically sharp predictions, allows us to elucidate the exact dependence of ICL performance on various hyperparameters and to pinpoint when and how the transition from memorization to ICL of linear regression occurs. In closing, we highlight the work of Reddy [21], who analyzed the emergence of in-context classification through a phenomenological model.
@@ II.2 Linear self-attention
where 0 in the lower-right corner is a token that prompts the missing value $y_{\ell+1}$ to be predicted.
@@
The output $A$ is a matrix while our goal is to predict a scalar, $y_{\ell+1}$. Following the choice of positional encoding in (6), we will take $A_{d+1,\ell+1}$, the element of $A$ corresponding to the 0 prompt, as the prediction for $y_{\ell+1}$:
@@ II.3 Pretraining data
[…]
@@
[…]
@@ II.4 Parameter reduction
where $\langle\cdot,\cdot\rangle$ stands for the standard inner product.
@@
We note that [29] provides an analysis of the population risk (whereas we focus on empirical risk) for a related reduced model in which both $v_{21}$ and $m_{21}$ are set to 0. Consequently, the predictors considered in [29] slightly differ from ours in (18)–(20) by an additive term (due to $m_{21}$). The authors of [29] justify this reduced model through an optimization argument: if these parameters $v_{21}$ and $m_{21}$ are initialized to zero, they remain zero under gradient descent optimization of the population risk.
@@ II.5 Model pretraining
$\Gamma^{\ast}$ $=\underset{\Gamma}{\arg\,\min}\,\sum_{\mu=1}^{n}\left(y_{\ell+1}^{\mu}-\langle\Gamma,H_{Z^{\mu}}\rangle\right)^{2}+\frac{n}{d}\lambda\mathinner{\!\left\lVert\Gamma\right\rVert}_{\mathrm{F}}^{2},$ (22)
[…]
$\text{vec}(\Gamma^{\ast})=\left(\frac{n}{d}\lambda I+\sum_{\mu=1}^{n}\text{vec}(H_{Z^{\mu}})\text{vec}(H_{Z^{\mu}})^{\top}\right)^{-1}\sum_{\mu=1}^{n}y_{\ell+1}^{\mu}\text{vec}(H_{Z^{\mu}}),$ (23)
@@ III.2 Learning curves for ICL of linear regression by a linear attention module
where $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ and $e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)$ are two deterministic functions of the parameters $\tau$, $\alpha$, $\kappa$, $\rho$ and $\lambda$. The exact expressions of these two functions can be found in Section E of the SI. For simplicity, we only present in what follows the ridgeless limit (i.e., $\lambda\to 0^{+}$) of the asymptotic generalization errors.
@@
where $\mathcal{M}_{\kappa}(\cdot)$, defined in Section H of the supplementary, is a function related to the Stieltjes transform of the Marchenko-Pastur law. Then
@@
$e^{\mathrm{IDG}}_{\mathrm{ridgeless}}\coloneqq\lim_{\lambda\to 0^{+}}e^{\mathrm{IDG}}(\tau,\alpha,\kappa,\rho,\lambda)=$ (34)
$\begin{cases}\frac{\tau}{1-\tau}\left(\frac{\rho+q^{\ast}-2q^{\ast}(1-\tau)({q^{\ast}}/{\xi^{\ast}}+1)}{1-p^{\ast}(1-\tau)}+\frac{\tau\mu^{\ast}(q^{\ast}+\xi^{\ast})^{2}}{q^{\ast}}\right)&\tau<1\\ \frac{\tau}{\tau-1}[\rho+q^{\ast}(1-q^{\ast}m^{\ast})]&\tau>1\end{cases},$ (35)
@@
We will discuss various implications of these equations in the next sections.
We derived these results using techniques from random matrix theory. The full setup and technical details are presented in the Supplementary Information. The computations involve analysis of the properties of the finite-sample optimal parameter matrix $\Gamma^{*}$ (see (23)). A key technical component of our analysis involves characterizing the spectral properties of the sample covariance matrix of $n=\Theta(d^{2})$ i.i.d. random vectors in dimension $\Theta(d^{2})$. Each of these vectors is constructed as the vectorized version of the matrix in (19). Related but simpler versions of this type of random matrices involving the tensor product of i.i.d. random vectors have been studied in recent work [39]. Some of our derivations are based on non-rigorous yet technically plausible heuristics. We support these predictions with numerical simulations and discuss in the Supplementary Information the steps required to achieve a fully rigorous proof.
@@ III.3 Sample-wise double-descent
Figure 1: ICL performance as a function of $\tau$: theory (solid lines) vs simulations (dots). Plots of (a.) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$, (b.) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$, and (c.) $e^{\mathrm{ICL}}(\tau,\alpha,\kappa,\rho,\lambda)$ against $\tau$. Simulated errors calculated by evaluating the corresponding test error on the corresponding optimized $\Gamma^{*}$. Parameters: $d=100$, $\rho=0.01$ for all; (a.), (b.) $\kappa=0.5$, (c.) $\alpha=10,\kappa=\infty$. Averages and standard deviations are computed over 10 runs.
How large should $n$, the pretraining dataset size, be for the linear attention to succesfully learn the task in-context? In Figure˜1, we plot our theoretical predictions for the ICL and IDG error as a function of $\tau=n/d^{2}$ and verify them with numerical simulations. Our results demonstrate that the quadratic scaling of sample size with input dimensions is indeed an appropriate regime where nontrivial learning phenomena can be observed.
As apparent in Figure˜1, we find that the generalization error for both ICL and IDG tasks are not monotonic in the number of samples. In the ridgeless limit, both ICL and IDG errors diverge at $\tau=1$, with the leading order behavior in the $\tau\uparrow 1$ (respectively $\tau\downarrow 1$) limit given by $\frac{c_{1}}{1-\tau}$ (respectively $\frac{c_{2}}{\tau-1}$), where $c_{1}$ (respectively $c_{2}$) is a $\tau$-independent constant. This leads to a “double-descent” behavior [40, 38] in the number of samples. As in other models exhibiting double-descent [40, 38, 41], the location of the divergence is at the interpolation threshold: the number of parameters of the model (elements of $\Gamma$) is, to leading order in $d$, equal to $d^{2}$, which matches the number of pretraining samples at $\tau=1$. Further, we can investigate the effect of ridge regularization on the steepness of the double descent, as illustrated in Figure˜1(c.) for the ICL task. As we would expect from other models exhibiting double-descent [40, 38, 41], increasing the regularization strength suppresses the peak in error around the interpolation threshold.
Finally, we note that if we take the limit of $\kappa\to\infty$ and $\alpha\to\infty$ in Result˜1 (in either order), the ICL generalization error in the ridgeless case reduces to the generalization error of simple ridgeless interpolation with isotropic Gaussian covariates in $d^{2}$ dimensions [38, 41]:
$\lim_{\alpha\to\infty}\lim_{\kappa\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}=\lim_{\kappa\to\infty}\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}=\begin{dcases}1-\tau+\frac{\rho}{1-\tau}&\tau<1,\\ \frac{\rho\tau}{\tau-1}&\tau>1.\end{dcases}$ (36)
@@ III.4 ICL and IDG error curves can have non-monotonic dependence on context length
Figure 2: Error curves as functions of $\alpha$: theory (solid lines) vs simulations (dots). Plots of (a.) $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ and (b.), (c.) $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ against $\alpha$. Hollowed markers, when plotted, indicate $\alpha^{\ast}$ value minimizing error if it exists. Parameters: $d=100$; (a.) $\tau=0.5,\rho=0.5$; (b.) $\tau=0.5,\rho=0.1$; (c.) $\tau=20,\rho=0.5$. Averages and standard deviations are computed over 20-100 runs.
How large should the context length be? In Figure˜2, we plot our theoretical results verified with experiments. We observe that we have correctly identified the regime where in-context and in-weight learning appear: context length indeed scales linearly with input dimensions, as numerical simulations computed using finite $d$ fit the asymptotic error curves.
@@
Figure 3: Phase diagram of non-monotonicity in $\alpha$ of ridgeless IDG and ICL error, for fixed $\tau$ against task diversity $\kappa$ and label noise $\rho$. The colormap corresponds to the value $\alpha^{\ast}$ that minimizes error in $\alpha$, if it exists, at that particular $\kappa,\rho$ pair; grey is plotted if the error curve is monotonic at that $\kappa,\rho$ pair. Figures (a.), (b.), (c.) in this plot correspond to the respective setup in Figure 2, and points A-E correspond to the respective curve in Figure 2. The dashed vertical lines are plotted at $\kappa=\text{min}(\tau,1)$. We know from (37) and surrounding discussion that ICL error diverges in $\alpha$ for all $\kappa,\rho$ to the left of this line, and IDG error diverges on this line for $\tau<1$.
While the functional form of $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}$ and $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ is too complex to study their minimizers analytically, we can investigate the non-monotonicity of IDG and ICL error in $\alpha$ numerically. This is done in Figure˜3, where the blue-green colormap corresponds to the value $\alpha^{\ast}$, if it exists, that minimizes $e^{\mathrm{IDG}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$ or $e^{\mathrm{ICL}}_{\mathrm{ridgeless}}(\tau,\alpha,\kappa,\rho)$ at fixed $\tau,\kappa,\rho$. The grey color indicates cases where the corresponding function is monotonic in $\alpha$. The results reveal that the non-monotonicity of the errors with respect to context length $\alpha$ depends in a nontrivial manner on both the task diversity $\kappa$ and the noise-to-signal ratio $\rho$.
@@
$\lim_{\alpha\to\infty}e^{\mathrm{ICL}}_{\mathrm{ridgeless}}$ $=\begin{dcases}\infty&\kappa\leq\text{min}(\tau,1),\\ 1-\tau+\rho+\frac{\rho\kappa\tau}{(\kappa-\tau)(1-\tau)}&\tau<1,\kappa>\tau,\\ \rho+\frac{\rho\kappa}{(\kappa-1)(\tau-1)}&\tau>1,\kappa>1.\end{dcases}$ (37)
Note the explicit divergence of ICL error for $\kappa\leq\text{min}(\tau,1)$ (independent of $\rho$). Similar investigation of Result 2 under the limit $\alpha\to\infty$ reveals an explicit divergence of IDG error for $\kappa=\tau$. The divergence in $\alpha$ of ICL error is a peculiarity of the ridgeless limit of errors we have taken in Results 1 and 2, and is further discussed in Section E of the Supplementary Information.